OpenStudy (kirbykirby):

For any 2 vectors a,b the cross product "(3a + 2b) x a" is the same vector as: a) 2a x b b) 2b x a c) 3a x b d) 3b x a e) 5b x a

OpenStudy (amistre64):

is cross product the same as dot product?

OpenStudy (amistre64):

i see that they arent.... i was sooo close lol

OpenStudy (kirbykirby):

yeah they aren't the same

OpenStudy (amistre64):

is cross priduct orthogonal? they perpendicular to each other then?

OpenStudy (kirbykirby):

yeah

OpenStudy (amistre64):

i was just reading up on vectors last night... if 2 vectors are othogonal, they have -1 as the product of their slopes... we wouldnt happen to know the vector quantities would we...

OpenStudy (kirbykirby):

no, the question I asked is really all the information given :(

OpenStudy (amistre64):

||a x b ||^2 = ||a||^2 ||b||^2 - (a.b)^2...does that help us any? :)

OpenStudy (kirbykirby):

is that like the law of cosines? o_O

OpenStudy (amistre64):

3<xa,ya> + 2<xb,yb> <3xa+2xb , 3ya + 2yb> right? lol.....wish I knew

OpenStudy (amistre64):

would it be asking this then? <3xa+2xb , 3ya + 2yb> (x) <xa,yx>

OpenStudy (kirbykirby):

oh perhaps

OpenStudy (kirbykirby):

well actually i think i'd need to add the z-component, since we'd get a 3-D vector

OpenStudy (amistre64):

i gots no idea how to add a z component in there ;)

OpenStudy (kirbykirby):

what would 3a x a = ? is it still a vector a multiplied by smtg?

OpenStudy (kirbykirby):

oh wait it would be 0 right o_o

OpenStudy (amistre64):

cant recall :) do you know where this mystery vector is supposed to be attached? I got this so far...but I may be wrong.

OpenStudy (kirbykirby):

ohh I figured it out! :) haha. Ok i wasn't sure but there's the property that (3a + 2b) x a = 3a x a + 2b x a = 0 + 2b x a... omg -_-

OpenStudy (amistre64):

thats a nice property to know then ;) good job!

OpenStudy (kirbykirby):

oye our prof never showed it to us :( Lol glad that was actually easy now xD thx for ur help nevertheless :)